OpenStudy (anonymous):

The volume of a cone is 300 cubic kilometers. If the radius of the base is 9 kilometers, what is the height of the cone? Use 3.14 for pi. show your work & round to the nearest 10th

OpenStudy (cwrw238):

Volume = (1/3)pi r^2 h

OpenStudy (cwrw238):

so height h = V / [(1/3)pir^2] so plug in pi=3.14, V = 300 , r = 9 and calculate h

OpenStudy (nurali):

volume V =1/3π(r^2)(h) so plug in pi=3.14, V = 300 , r = 9 h = 300×3/(36π) = 7.96178 = 8.0 km